OpenStudy (anonymous):

determine whether (-1,-2) is a solution of 4x+9y=-6. I just need the answers please. I have 28 questions I need to get thru real quick. Thanks

OpenStudy (sandra):

4*-1 + 9*-2 = -4 -18 = -22

OpenStudy (sandra):

-22 != -6

OpenStudy (sandra):

so no